### v1

# This work has been submitted to an IEEE conference for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. Mobile Jamming Mitigation in 5G Networks: A MUSIC-Based Adaptive Beamforming Approach

Abstract
Mobile jammers pose a critical threat to 5G networks, particularly in military communications. We propose an intelligent anti-jamming framework that integrates Multiple Signal Classification (MUSIC) for high-resolution Direction-of-Arrival (DoA) estimation, Minimum Variance Distortionless Response (MVDR) beamforming for adaptive interference suppression, and machine learning (ML) to enhance DoA prediction for mobile jammers. Extensive simulations in a realistic highway scenario demonstrate that our hybrid approach achieves an average Signal-to-Noise Ratio (SNR) improvement of 9.58 dB (maximum 11.08 dB) and up to 99.8% DoA estimation accuracy. The framework’s computational efficiency and adaptability to dynamic jammer mobility patterns outperform conventional anti-jamming techniques, making it a robust solution for securing 5G communications in contested environments.
Index Terms:
Anti-Jamming, Adaptive Beamforming, 5G, Multiple Signal Classification, Minimum Variance Distortionless Response, Barrage Jamming, Direction of Arrival
I Introduction
The widespread adoption of 5G networks has transformed wireless communications, enabling high-speed and low-latency connectivity for a range of critical applications. However, these networks are increasingly vulnerable to sophisticated jamming attacks, particularly from mobile jammers that dynamically change their position to disrupt reliable communication [1]. Beamforming, a fundamental technology in 5G, enhances signal quality and mitigates interference but faces significant challenges against mobile jammers, as traditional anti-jamming methods rely on static assumptions unsuitable for dynamic interference sources.
Conventional interference mitigation strategies, such as fixed beamforming patterns, are effective against stationary jammers but fail to adapt to the mobility of jammers, which degrade network performance by introducing intentional interference across the spectrum [1]. Recent advances in adaptive beamforming, particularly combining Multiple Signal Classification (MUSIC) for Direction-of-Arrival (DoA) estimation with Minimum Variance Distortionless Response (MVDR) for interference suppression, offer promising solutions. However, these approaches often lack the real-time adaptability needed to counter unpredictable jammer mobility patterns, a critical gap for ensuring robust 5G communications in dynamic environments [2].
Figure 1: Simulated highway scenario illustrating the interaction between a stationary transmitter (Tx), a mobile receiver (Rx), and a mobile jammer (Jammer), designed to evaluate dynamic anti-jamming performance.
This paper proposes an intelligent anti-jamming framework to address mobile jammers in 5G networks, leveraging adaptive beamforming optimized for dynamic scenarios, as shown in Fig. 1. Our approach integrates an enhanced MUSIC algorithm for high-resolution DoA estimation, MVDR beamforming for adaptive interference suppression, and machine learning (ML) to improve DoA prediction under varying jammer mobility. Key contributions include: (1) a novel MUSIC-ML hybrid for accurate DoA estimation, (2) tailored MVDR beamforming for mobile jammer mitigation, and (3) comprehensive performance analysis demonstrating significant Signal-to-Noise Ratio (SNR) improvements. The framework demonstrates robust performance in mobile jamming scenarios while maintaining computational efficiency.
The remainder of this paper is structured in the following way: Section II reviews background and related work. Section III describes the proposed MUSIC-MVDR-ML framework. Section IV presents experimental results, discussion, and limitations. Finally, Section V concludes the paper and outlines future research directions.
II Background and Related Work
Beamforming, essential for enhancing signal quality and mitigating interference in modern wireless communications like 5G, has evolved from conventional fixed-weight methods to adaptive solutions [3]. While conventional beamforming offers simplicity, its fixed patterns struggle against dynamic interference, particularly mobile jammers. Adaptive techniques like MVDR and Linear-Constrained Minimum Variance (LCMV) adjust parameters based on the signal environment.
MVDR beamforming optimizes array weights to maximize sensitivity towards a desired DoA while suppressing other sources [4]. However, its performance degrades if the desired DoA is inaccurate. To address this, MVDR is coupled with high-resolution DoA estimation algorithms like MUSIC, which uses eigenspace decomposition to accurately identify signal directions even in multi-source scenarios [5, 6]. LCMV also minimizes interference but imposes linear constraints to preserve the target signal [7]. A practical limitation is that LCMV is often restricted to a uniform linear array (ULA), whereas MVDR offers more flexibility, adapting well to a uniform rectangular array (URA) used in this work [6].
Traditional anti-jamming strategies focusing on spatial filtering are effective against stationary jammers but fail when confronting mobile ones that dynamically change positions [8]. At the network layer, multipath communications and secret sharing schemes enhance resilience by distributing data across multiple paths to counter jamming attacks [9, 10]. The combination of MUSIC’s accurate DoA estimation with ML support and MVDR’s adaptive nulling presents a promising approach for mobile interference mitigation.
Finally, accurately evaluating performance through signal-to-noise ratio (SNR) in multi-antenna systems facing mobile jammers presents challenges. Standard SNR calculations may not adequately handle dynamically changing interference or the specifics of signal combination after beamforming, necessitating careful consideration of the calculation methodology, as addressed in our proposed framework. The limitations of current DoA estimation with mobile sources, real-time adaptation challenges, and the need for better ML integration highlight the necessity for the comprehensive approach developed in this paper.
III Proposed Framework
Our proposed framework integrates MUSIC-based DoA estimation, MVDR beamforming, and ML enhancement for DoA estimation support. We consider a $2\times 2$ URA that receives signals from a desired transmitter and a mobile jammer. The received signal $x(t)$ at the array is modeled as:
$x(t)=s(t)a(\theta_{s},\phi_{s})+j(t)a(\theta_{j},\phi_{j})+n(t)$ (1)
where $s(t)$ is the desired signal, $j(t)$ is the jamming signal, $n(t)$ is additive white Gaussian noise, and $a(\theta,\phi)$ is the array steering vector which corresponds to azimuth angle $\theta$ and elevation angle $\phi$. Subscripts $s$ and $j$ denote the desired signal and jammer, respectively.
III-A Anti-Jamming System
MUSIC-Enhanced DoA Estimation: The MUSIC algorithm estimates an angle of arrival by exploiting the eigenstructure of the received signal’s covariance matrix $R$ [5]:
$R=E\{x(t)x^{H}(t)\}$ (2)
Eigendecomposition of $R$ separates the signal subspace ($U_{s}$) from the noise subspace ($U_{n}$) [11]:
$R=U_{s}\Lambda_{s}U_{s}^{H}+U_{n}\Lambda_{n}U_{n}^{H}$ (3)
This allows for high-resolution DoA estimation by using $R$ and scanning over a grid of possible angles, in our case -90 to 90, then evaluating the MUSIC pseudospectrum. Peaks in the pseudospectrum show the estimated DoA of an incoming signal.
Adaptive MVDR Beamforming: Beamforming is a spatial filtering strategy where various weight geometries and gain tapering methods manipulate transmitter and receiver antennas. A conventional beamformer is designed to stay the same between collections, while an adaptive beamformer responds to environment, allowing the antenna array to dynamically change. The MVDR beamformer is an adaptive beamformer that computes optimal weights $w_{opt}$ to minimize output power while preserving the desired signal gain (in the direction $\theta_{s}$, assumed to be the desired signal’s DoA) [12]:
$w_{opt}=\frac{R^{-1}a(\theta_{s},\phi_{s})}{a^{H}(\theta_{s},\phi_{s})R^{-1}a(\theta_{s},\phi_{s})}$ (4)
This process adaptively creates nulls in the direction of interference sources such as the mobile jammer. In order to effectively shape the beam, we apply $w_{opt}$ to the antenna array.
Machine Learning Integration: To enhance estimation accuracy, particularly when the initial MUSIC DoA estimation error exceeds a predefined threshold, a linear regression ML model is employed. Linear regression was chosen due to its simplicity in implementation and relationship between input output. The model is trained using MATLAB’s fitlm [13] where the input features (predictors) are sequences of three consecutive ground truth azimuth angle values, and the output (response) is the subsequent ground truth azimuth value. Over 500 DoA data points from the simulated mobility patterns of the scenario were used as training data shown in Fig. 2. The model predicts the DoA’s next azimuth angle $\hat{\theta}_{t+1}$ based on the past $k$ measurements $(\theta_{t},\theta_{t-1},...,\theta_{t-k})$:
$\hat{\theta}_{t+1}$ $=\beta_{0}+\sum_{i=1}^{k}\beta_{i}\cdot\theta_{t-i}$ (5)
While $k=3$ is only used in our experiments, the above linear prediction model can be easily generalized by leverging more historical measurements. The high statistical significance of the $\theta_{t-2}$ coefficient (p-value $\approx 3.4\times 10^{-76}$) in our experiments confirms strong temporal dependency, justifying the model’s predictive capability.
Figure 2: Optimized linear regression model showing relationship between the adjusted whole model values and the adjusted data with 95% confidence bounds.
III-B Performance Quantification
SNR Calculation: To reliably evaluate performance, we use a specific SNR calculation method that averages antenna signals across the $M\times N$ array elements to prevent artificial gain inflation and ensures consistency between simulation and potential implementation, both before and after MVDR processing. The combined signal is $x_{2}(t)=\frac{1}{N+M}\sum_{i=1}^{N+M}|x(t)_{i}|$. Signal and noise powers are calculated by averaging $|x_{2}(t)|$ over known signal-present ($a_{1}$ to $b_{1}$) and signal-absent ($a_{0}$ to $b_{0}$) regions:
Signal Power $=\frac{1}{b_{1}-a_{1}+1}\sum_{i=a_{1}}^{b_{1}}|x_{2}(i)|$ (6)
Noise Power $=\frac{1}{b_{0}-a_{0}+1}\sum_{i=a_{0}}^{b_{0}}|x_{2}(i)|$ (7)
SNR (dB) $=20\log_{10}\left(\frac{\text{Signal Power}}{\text{Noise Power}}\right)$ (8)
The SNR improvement is calculated as:
$\text{SNR}_{improvement}=\text{SNR}_{after}-\text{SNR}_{before}$ (9)
Figure 3: Simulation framework for mobile jamming mitigation with ML-enabled dynamic beamforming.
Mitigation Algorithm: Fig. 3 is a flow diagram of Algorithm 1 which outlines the overall simulation framework to mitigate jamming in mobile environment, which integrates MUSIC-based DoA estimation, MVDR beamforming, and ML prediction model. It iteratively collects signals, estimates DoA using MUSIC (potentially refined by the ML model when MUSIC error is high, as described above), computes MVDR weights, applies beamforming, and evaluates performance based on the SNR improvement. The system can adapt parameters or retrain the ML model, if necessary, to further improve the performance. This integrated approach provides robust protection against mobile jammers while maintaining computational efficiency.
Algorithm 1 Mobile Jammer Mitigation
1:Received signal stream $x(t)$, array parameters, ML model
2:Optimized MVDR weights, enhanced SNR performance
3:Initialize: Array parameters, ML model, and DoA threshold $\epsilon$
4:while System is operational do
5:     Collect current signals $x(t)$
6:     Compute covariance matrix $R\leftarrow E[x(t)x^{H}(t)]$
7:     Estimate DoA angles $\{\theta_{s},\,\phi_{s}\}\leftarrow\mathrm{MUSIC}(R)$
8:     if $DoA$ $Error>\epsilon$ then
9:         Predict $\theta_{s}$ as $\hat{\theta}_{t+1}$ using ML model
10:     end if
11:     Update ML model with latest measurements (e.g., recent DoAs for next prediction)
12:     Compute MVDR weights $w_{\mathrm{opt}}\leftarrow\frac{R^{-1}a(\theta_{s},\phi_{s})}{a^{H}(\theta_{s},\phi_{s})\,R^{-1}\,a(\theta_{s},\phi_{s})}$
13:     Apply beamforming with $w_{\mathrm{opt}}$ and measure $\text{SNR}_{{after}}$ using Eq. (8)
14:     Calculate $\text{SNR}_{improvement}$ using Eq. (9)
15:end while
IV Experiment
IV-A Setup
We evaluated the proposed framework through extensive MATLAB simulations emulating a realistic highway scenario as seen in Fig. 1. This scenario involved a stationary transmitter (cell tower), a mobile receiver (red car), and a mobile jammer (blue van) moving along various predefined paths. The receiver and jammer have the ability to begin in any of the three cardinal directions (where the receiver in the figure began on the left-hand side of the scenario, which is equivalent to west). Trajectory options include North-to-South, North-to-West, South-to-North, South-to-West, West-to-North, and West-to-South. Key simulation parameters include:
The carrier frequency of 100 MHz was selected to simplify simulation and reduce computational complexity, as the MUSIC-MVDR-ML framework is frequency-agnostic and can be adapted to standard 5G bands like 3.5 GHz in future implementations. Similarly, the signal sampling rate of 1 kHz was chosen for baseband signal processing after downconversion, which is sufficient for the simulated scenario.
  • Jammer: Mobile or Stationary
  • Carrier frequency: 100 MHz
  • Signal sampling rate: 1 kHz
  • Antenna array: $2\times 2$ URA
  • Jammer power: 10 dB
IV-B Results
DoA Estimation and ML Performance: The integrated MUSIC algorithm demonstrated high DoA estimation accuracy (up to 99.8%) with a dynamic jammer present across the tested scenarios. Fig. 4 shows an example of a beam pattern with the estimated DoA indicated by its peak at -40. The machine learning support aids in this estimation when error is too high.
Adaptive Beamforming: The adaptive MVDR beamformer effectively steered nulls towards the jammer direction while preserving the desired signal. Analysis of the beam patterns (e.g., Fig. 4) confirmed improved directionality and interference suppression compared to initial states. The system successfully mitigated jamming interference along both linear and diagonal trajectories within the highway environment.
SNR Enhancement: Using our defined SNR calculation method, which averages the combined antenna signals, the framework achieved significant improvement. Across the six diverse highway trajectories simulated, the system yielded an average SNR improvement of 9.58 dB post-MVDR beamforming. While the improvement varied depending on the relative positions and path (as illustrated in Fig. 5), for all trajectories there is a spike in SNR at around sample 25 that corresponds to the intersection point of the simulated highway. The enhancement was consistently observed, validating the system’s anti-jamming capability.
Figure 4: Example MVDR beam pattern demonstrating effective null steering towards the jammer direction (e.g., at 30) while maintaining gain towards the desired signal (e.g., at -40).
Computational Efficiency: The framework demonstrated suitability for real-time application, with average processing times per collection measured as: DoA estimation (MUSIC) $\approx$ 41.9 ms, MVDR weight computation $\approx$ 5.2 ms, and ML prediction $\approx$ 16.9 ms111github.com/CLIS-WPI/mobile-jamming-mitigation..
Baseline Comparison: As previously mentioned, adaptive beamforming is used in this anti-jamming system as well as machine learning. The following points highlight direct comparisons among different MATLAB simulations. The values indicate SNR improvement, DoA accuracy, and processing time across three methods:
  • MUSIC-MVDR-ML (Proposed): 9.58 dB, 99.8%, 63.0 ms
  • MUSIC-MVDR: 9.46 dB, 99.3%, 74.4 ms
  • Fixed Beamforming: 1.90 dB, 88.0%, 240.2 ms
IV-C Discussion
Limitations: Despite robust performance, limitations exist in this adaptive beamforming for anti-jamming setup. SNR calculation accuracy can decrease in very high-mobility scenarios. The current ML model might require retraining for drastically different mobility patterns not encountered during training. Furthermore, the framework’s performance against multiple simultaneous jammers needs investigation. Only one jammer was implemented for sake of simplicity. To introduce a multi-jammer system more jammer objects and channels would need to be initalized and added at collection time. Lastly, due to the nature of the simulation, the propagation of signals occurs at every collection prior to the processing step inside the iterative for loop. This causes latency within the system, as multi-threading was not explored.
Figure 5: SNR improvement achieved by the proposed MUSIC-MVDR-ML framework along a representative highway trajectory, showcasing consistent enhancement despite jammer mobility.
TABLE I: Summary of Key Performance Metrics
Metric Value
Average SNR Improvement 9.58 dB
Maximum SNR Improvement 11.08 dB
DoA Estimation Accuracy Up to 99.8%
Avg. Processing Time per Collection:
   - MUSIC $\approx$ 41.9 ms
   - MVDR $\approx$ 5.2 ms
   - ML $\approx$ 16.9 ms
V Conclusion and Future Work
This paper presented a novel anti-jamming framework integrating MUSIC DoA estimation, MVDR beamforming, and machine learning to mitigate mobile 5G jammers. Simulations confirmed its effectiveness, achieving an average 9.58 dB SNR improvement and high DoA accuracy (up to 99.8%) in realistic highway scenarios. Future work will focus on addressing the limitations, potentially exploring deep learning models and optimizing for multi-jammer environments. This project demonstrated adaptive beamforming for anti-jamming as a proof of concept in simulation. Next steps would be implementation in a lab setting with USRP software defined radios and PCB directional antennas.
Acknowledgment
This material is based upon work supported in part by NSF under Awards CNS-2120442 and IIS-2325863, and NTIA under Award No. 51-60-IF007. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the NSF and NTIA.